OpenStudy (anonymous):

write in logarithmic form (11/2)^4 = 16/14641

OpenStudy (shadowfiend):

Remember, if: \[a^b = c\] Then: \[log_a c = b\] So can you see how you could rewrite your expression in logarithmic form?

OpenStudy (anonymous):

log 11/2 16/14641 = 4 ?

OpenStudy (shadowfiend):

Right, base 11/2, so: \[\log_{11/2} \frac{16}{14641} = 4\]

OpenStudy (shadowfiend):

Depending on how much you know about logs and such, you can simplify that further to: \[\log_{11/2}16 - \log_{11/2}14641 = 4\]

OpenStudy (anonymous):

thanks

OpenStudy (shadowfiend):

No problem, glad to help :)